OpenStudy (anonymous):

find the inverse of the given function. f(x)=3/24x-13

OpenStudy (anonymous):

Find x in terms of y.. Firstly let f(x) = y \[y = \frac{3x}{24} - 13\] \[24y = 3x - 13*24\] \[3x = 24y + 13 \times 24\] solve this to get x.. That will be the inverse..

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Mention not..